### v1

# White-Basilisk: A Hybrid Model for Code Vulnerability Detection

Abstract
The proliferation of software vulnerabilities presents a significant challenge to cybersecurity, necessitating more effective detection methodologies. We introduce White-Basilisk, a novel approach to vulnerability detection that demonstrates superior performance while challenging prevailing assumptions in AI model scaling. Utilizing an innovative architecture that integrates Mamba layers, linear self-attention, and a Mixture of Experts framework, White-Basilisk achieves state-of-the-art results in vulnerability detection tasks with a parameter count of only 200M. The model’s capacity to process sequences of unprecedented length enables comprehensive analysis of extensive codebases in a single pass, surpassing the context limitations of current Large Language Models (LLMs). White-Basilisk exhibits robust performance on imbalanced, real-world datasets, while maintaining computational efficiency that facilitates deployment across diverse organizational scales. This research not only establishes new benchmarks in code security but also provides empirical evidence that compact, efficiently designed models can outperform larger counterparts in specialized tasks, potentially redefining optimization strategies in AI development for domain-specific applications.
1 Introduction
Recent advancements in artificial intelligence have been largely driven by the rapid expansion of neural networks, where cutting-edge performance is often linked to models that scale to hundreds of billions of parameters. While this scaling paradigm has yielded impressive capabilities, it has also introduced significant computational, environmental, and accessibility challenges (Patterson et al., 2021; Faiz et al., 2023). As the field continues to prioritize ever-larger models, an important question emerges: are there domains where architectural innovation and targeted design might outperform sheer parameter count? This question becomes particularly relevant for high-stakes applications with specific computational constraints, where the "bigger is better" approach may prove suboptimal or economically infeasible.
Code vulnerability detection represents precisely such a domain—one with substantial practical importance and unique technical challenges. The economic implications are significant: cybercrime is projected to reach \$10.5 trillion by 2025 (Cybersecurity Ventures, 2023). This massive economic impact is largely preventable, as the Ponemon Institute found that 60% of breach victims were compromised due to an unpatched known vulnerability, while 62% of organizations reported they were completely unaware that their systems were vulnerable prior to a data breach (Ponemon Institute, 2018). When vulnerabilities are found in third-party software, the financial consequences are severe, costing organizations an average of \$4.55 million per breach incident (UpGuard, 2023). Despite this growing threat landscape and expanding attack surface resulting from increasingly complex software systems, existing automated vulnerability detection approaches face significant limitations. Traditional static application security testing (SAST) tools demonstrate insufficient detection rates (Zhou et al., 2024), while recent machine learning approaches based on large language models (LLMs) require prohibitive computational resources, making them impractical for many real-world deployment scenarios. Moreover, detecting vulnerabilities often requires analysing extremely long code sequences and capturing subtle interactions spanning multiple functions or files—capabilities that remain challenging even for the largest models due to context length constraints and inefficient attention mechanisms.
Current research has explored various approaches to address these challenges. Transformer-based models fine-tuned on code have shown promising results (Feng et al., 2020; Guo et al., 2022), but their quadratic attention complexity limits their application to large codebases. Specialized vulnerability detection architectures (Hanif and Maffeis, 2022; Fu and Tantithamthavorn, 2022) improve upon these foundations but still struggle with long-range dependencies and resource efficiency. Notably, even as these models grow in size and complexity, their performance on realistic vulnerability detection benchmarks often falls short of practical requirements (Ding et al., 2024). This disconnect between model scale and real-world effectiveness suggests an opportunity to re-examine fundamental architectural assumptions in the context of this specialized task.
We introduce White-Basilisk, a compact 200M-parameter model that challenges conventional scaling wisdom by outperforming substantially larger models on code vulnerability detection tasks. Our approach diverges from the standard practice of scaling existing architectures, instead focusing on designing a specialized hybrid architecture that addresses the unique requirements of vulnerability detection. The key insight driving our work is that by combining state-space models (Mamba) with linear-complexity attention mechanisms and sparsely-activated Mixture of Experts, we can efficiently process extremely long code sequences while capturing both local syntactic patterns and global dependencies crucial for vulnerability identification.
Our contributions can be summarized as follows:
  1. 1.
    Novel Hybrid Architecture: We propose a resource-efficient model integrating Mamba layers for local pattern recognition, linear-complexity attention for global context modeling, and conditional computation through Mixture of Experts, enabling effective vulnerability detection with just 200M parameters.
  2. 2.
    Extended Context Processing: White-Basilisk analyses sequences up to 128,000 tokens during inference—over 8× longer than typical transformer models—enabling comprehensive analysis of entire codebases on a single GPU and detection of vulnerabilities spanning multiple functions or files.
  3. 3.
    State-of-the-Art Performance: Through rigorous evaluation across five established benchmarks (PRIMEVUL Ding et al. (2024), BigVul Fan et al. (2020), Draper Russell et al. (2018), REVEAL Chakraborty et al. (2021), and VulDeepecker Li et al. (2018)), we demonstrate that White-Basilisk outperforms models up to 35× larger, particularly on realistic, imbalanced datasets representative of real-world code security scenarios.
  4. 4.
    Environmental and Computational Efficiency: We provide a detailed analysis of White-Basilisk’s carbon footprint and computational requirements, demonstrating up to 99.9% reduction in emissions compared to larger models while maintaining superior performance—establishing a new efficiency frontier for secure, sustainable AI systems.
Beyond the immediate applications in software security, our work presents broader implications for the field of AI. By demonstrating that a carefully designed domain-specific architecture can outperform general-purpose scaled models, we challenge the notion that parameter count is the primary determinant of model capability. White-Basilisk serves as evidence that the path toward more capable AI systems may lie not only in scaling existing architectures but also in rethinking fundamental design choices for specific high-value applications. In the following sections, we detail White-Basilisk’s architecture, training methodology, and comprehensive evaluation across multiple vulnerability detection benchmarks.
2 Related Work
The field of automated code vulnerability detection provides an excellent context for exploring efficient AI architectures. While this domain has undergone rapid evolution, most approaches have followed the general AI trend toward larger models with increasing computational requirements. Our work challenges this paradigm by demonstrating that carefully designed smaller models can achieve superior results. To properly position our contribution, we review both traditional vulnerability detection approaches and recent trends in efficient language modeling.
Driven by the growing complexity and security risks of software systems, early pioneers explored static analysis techniques Chess and McGraw (2004) and pattern matching methods Livshits and Lam (2005). While these approaches laid a foundational framework, they often encountered significant drawbacks, including high false positive rates, difficulty detecting complex vulnerabilities, and susceptibility to obfuscation techniques. Due to these limitations, they were generally unsuitable for active production environments.
As the field matured, researchers recognized the transformative potential of AI, gradually shifting from conventional practices to a new paradigm. This transition was marked by the development of VulDeePecker Li et al. (2018), one of the first deep learning (DL)-based systems for vulnerability detection. VulDeePecker utilized code gadgets and Bidirectional Long Short-Term Memory (BiLSTM) networks to identify vulnerabilities in C/C++ code. This work demonstrated the ability of DL techniques to capture complex patterns associated with code vulnerabilities, paving the way for the development of further ML-driven solutions. However, its reliance on manually crafted features limited its generalizability. Building on this work, Russell et al. (2018) developed the Draper dataset, which provides a substantial real-world dataset specifically designed for training neural networks in the task of vulnerability detection. Their work showed the advantages of leveraging vast training data to enhance model performance, improving performance but still struggling with limited context windows that restricted the capture of long-range dependencies in code.
Following that, pre-trained LLMs emerged as a significant breakthrough in code analysis. Hanif and Maffeis (2022) proposed VulBERTa, an adaptation of the RoBERTa model for detecting code vulnerabilities, demonstrating the potential of transfer learning from natural language processing to code analysis. By fine-tuning LLMs pre-trained on extensive corpora of code, this approach quickly gained popularity due to its ability to capture latent patterns in both code structure and semantics. However, similar to many transformer-based models, it suffers from quadratic computational complexity with sequence length, constraining its applicability to large-scale projects.
Another line of research focused on models specifically tailored for code understanding. Feng et al. (2020) introduced CodeBERT, a bimodal pre-trained model for programming language and natural language. This work allowed for a deeper understanding of both code and its associated documentation, proving useful across a range of software engineering tasks, including vulnerability detection. However, while powerful, these models typically demand significant computational resources and encounter difficulties when processing extremely long sequences, which restricts their practical use in analyzing large codebases. For instance, CodeBERT was pre-trained for approximately 1000 hours on 16 interconnected NVIDIA Tesla V100 GPUs, representing a substantial energy and resource investment.
Other recent work concentrated on improving the granularity and efficiency of vulnerability detection. Such an example is LineVul (Fu and Tantithamthavorn (2022)), a transformer-based approach for line-level vulnerability prediction, enabling precise localization of vulnerabilities within codebases. While valuable for precise localization, this approach may overlook vulnerabilities that span multiple lines or functions.
Comprehensive studies by Chakraborty et al. (2021) and Ding et al. (2024) highlighted persistent challenges in the field, including data quality issues, unrealistic evaluation methods, and difficulties in handling long-range dependencies. Ding et al. (2024) showed that existing benchmarks significantly overestimate model performance, with state-of-the-art models achieving high scores on flawed datasets but failing on more realistic ones. Our work directly addresses these concerns through improved data collection, realistic evaluation metrics, and model architecture designed for long-range understanding.
Currently, existing vulnerability detection methods primarily focus on their main objectives without considering the increasing size of model parameters. This uncontrolled growth in model size has led to the development of energy-hungry models requiring dozens of cutting-edge GPUs. Most recently, research developments have been made in efficient language model design. For example, building on the work of Mamba Gu and Dao (2023), Lieber et al. (2024) introduced Jamba, which employs a Mixture of Experts (MoE) strategy, combining Mamba layers and attention mechanisms in an interleaving pattern Fedus et al. (2022). In addition, Wang et al. (2020) proposed the Linformer, which can reduce self-attention complexity to linear time using low-rank approximations, while the Transformer-XL architecture Dai et al. (2019) has been shown to effectively model long-term dependencies by introducing recurrence into self-attention.
3 Model Architecture
The architecture of White-Basilisk was designed to address three key challenges in code vulnerability detection. Firstly, it tackles the problem of long-range dependencies, as code vulnerabilities often span multiple functions or even files, necessitating a model capable of understanding extensive contexts. Secondly, the architecture strikes a balance between local and global information processing. This dual focus enables both fine-grained understanding of code syntax and broad comprehension of overall program structure, both crucial for effective vulnerability detection. Lastly, the design prioritizes computational efficiency, aiming to create a powerful model that can be deployed in real-world settings without requiring massive computational resources. By addressing these interconnected objectives, our model provides a robust solution for identifying and analyzing potential security flaws in code. To this end, we developed a novel hybrid architecture that combines three main components:
  1. 1.
    Mamba layers: These form the backbone of our model, efficiently capturing local dependencies and providing adaptive computation based on input content.
  2. 2.
    White-Basilisk Self-Attention: Our novel adaptation of linear attention mechanisms enables efficient processing of extremely long sequences, enabling whole-codebase analysis.
  3. 3.
    Mixture of Experts (MoE): This adds dynamic adaptability throughout the network, with MoE layers configured with 8 Experts and 2 Experts per token.
The synergy between these components allows White-Basilisk to process sequences up to 128,000 tokens during inference, a capability that sets it apart in the field of code analysis. This extensive context window enables holistic analysis of entire codebases, potentially uncovering vulnerabilities that span multiple functions or files.
Mamba layers form the backbone of White-Basilisk, chosen for their exceptional efficiency in capturing local dependencies in code sequences. Unlike traditional recurrent neural networks (RNNs) or attention mechanisms, Mamba’s selective state-space mechanism allows for linear-time computation with respect to sequence length. In White-Basilisk, Mamba layers are implemented with a State size ($d_{\text{state}}$) of 16, a Convolution kernel size ($d_{\text{conv}}$) of 4 and an Expansion factor of 2.
The Mixture of Experts (MoE) layers in White-Basilisk introduce dynamic adaptability while maintaining computational efficiency. Configured with 8 Experts and 2 Experts per token, these layers allow the model to activate only a subset of parameters for each input token, significantly reducing computational cost compared to fully-dense models of similar capacity. This sparsity contributes to faster inference times, a critical factor in real-world deployment scenarios.
The layer combination pipeline, inspired by the Jamba model Lieber et al. (2024), allows for a sophisticated interleaving pattern. Specifically, the layer combination is defined by two main configuration parameters determined through experimentation: the attention layer offset (2) and the attention layer period (8). The conjunction of layers in our architecture is defined by the following formula:
$\text{Layer}_{i}=\begin{cases}\text{Attention}(x),&\text{if }(i-2)\bmod 8=0\text{ and }i\geq 2\\ \text{MoE}(x),&\text{if }i\bmod 2=1\\ \text{Mamba}(x),&\text{otherwise}\end{cases}$ (1)
The forward pass through the model can be expressed as $h_{i}$, where $i\in\{0,1,...,L-1\}$ is the layer index and $L$ is the total number of layers in the architecture. Here, $h_{i}$ is the hidden state after the $i$-th layer, and $h_{0}$ is the initial input embedding. The final output of the model $y$ is obtained by applying layer normalization to the last hidden state $h_{L}$:
$h_{i}=\text{Layer}_{i}(h_{i-1})+h_{i-1}\quad\text{and}\quad y=\text{LayerNorm}(h_{L})$ (2)
The residual connections facilitate gradient flow during training and allow for the preservation of information across layers.
3.1 White-Basilisk Self-Attention: Our Novel Linear-Scaling Mechanism
A key innovation in White-Basilisk is our novel self-attention mechanism that achieves linear complexity with respect to sequence length, enabling the model to efficiently process entire codebases up to 128K tokens. Traditional attention mechanisms face a fundamental scaling limitation due to their quadratic complexity, making them computationally prohibitive for analyzing large codebases. Our approach transforms this limitation while maintaining attention quality.
Taking inspiration from recent advances in efficient attention Munkhdalai et al. (2024), we developed a substantially modified implementation that prioritizes holistic sequence understanding rather than segmented processing. The primary distinctions in our approach are:
  1. 1.
    Global Context Accumulation: Unlike segmented approaches that process chunks independently, White-Basilisk Self-Attention accumulates context information across the entire sequence:
    $\text{total}_{\text{mem}}=\sum_{s=1}^{S}A_{\text{mem,s}},\quad\text{total}_{\text{attn}}=\sum_{s=1}^{S}A_{\text{dot,s}}$ (3)
    where $S$ is the total number of segments. This accumulation leads to linear memory growth with sequence length while enabling attention across the full context.
  2. 2.
    Adaptive Global Gating: Our approach implements a novel global gating mechanism that operates on the entire accumulated context:
    $O=\text{sigmoid}(\beta)\odot\text{total}_{\text{mem}}+(1-\text{sigmoid}(\beta))\odot\text{total}_{\text{attn}}$ (4)
    This allows for adaptive balancing of memory-based and direct attention components across the entire sequence, unlike traditional approaches that gate each segment individually.
The memory-based component of our attention mechanism adaptively compresses and retrieves context information through:
$A_{\text{mem}}=\frac{(\text{ELU}(Q)+1)M^{T}}{(\text{ELU}(Q)+1)z^{T}+\epsilon}$ (5)
$M_{\text{new}}=M+(\text{ELU}(K)^{T}+1)V,\quad z_{\text{new}}=z+\sum_{i=1}^{L}(\text{ELU}(K_{i})+1)$ (6)
where $M$ is the compressive memory, $z$ is the normalization term, and $L$ is the segment length.
A critical aspect of our implementation is the use of learnable parameters to optimize both the memory accumulation and attention distribution. The memory matrix $M$ incorporates learnable projection matrices that are trained to retain information that might be missed by the local processing of Mamba layers. Through shared residual connections with surrounding Mamba layers (as expressed in the forward pass equations), our attention mechanism effectively learns to complement the local processing strengths of Mamba with global context awareness.
This integration is particularly notable in how it differs from the original Infini-attention approach, which processes segments independently. Our adaptation maintains continuous information flow across the entire sequence through the accumulation of memory states, allowing the model to establish connections between distantly separated code elements. This capability is essential for detecting vulnerabilities that span multiple functions—for example, identifying a buffer allocation in one function and its potential overflow in another much later in the codebase.
4 Experimental Setup
The development and evaluation of source code vulnerability detection models requires a large collection of annotated data samples. In this section, we outline the datasets chosen for this purpose and explain how they were used for both model training and testing purposes. Additionally, we provide a detailed overview of the training methodology used for our model.
4.1 Pretraining
Traditional LLM training methods are generally designed to enable models to comprehend language and its syntax. This is often accomplished through Causal Language Modeling (CLM), where a model learns to predict the next token given its input. Another common approach is based on the Fill in the Middle (FIM) technique, in which random text portions are masked, and the model must reconstruct the missing content. Some advanced source code LLMs combine both methods to increase model flexibility (Li et al., 2023). Similarly, in our work, we employ both techniques during model pre-training on the selected 2M code samples. This process requires approximately 600 hours to complete on a single NVIDIA A100 40GB GPU.
For pretraining data, we initially selected a carefully curated subset of the StarCoder dataset (Li et al., 2023), which includes more than 80 programming languages and consists of 305M files in total. For our study, we focused on C and C++ code samples, using 2M code samples during the pre-training phase. This focused approach allowed us to achieve excellent general performance while maintaining a relatively small pretraining dataset, challenging the notion that massive pretraining datasets are always necessary for high-performance models.
4.2 Fine-tuning and Evaluation
To evaluate the pre-trained model, we required well-established benchmarking datasets with publicly available partitions for fine-tuning and testing. This ensures a fair comparison with existing methods without the need to recreate the original models. For this purpose, we selected five publicly available datasets: VulDeePecker (Li et al., 2018), Draper (Russell et al., 2018), PrimeVul (Ding et al., 2024), REVEAL (Chakraborty et al., 2021), and BigVul (Fan et al., 2020).
These datasets present significant challenges that validate White-Basilisk’s capabilities. They feature significant class imbalance with the vulnerable class representing only 2.9-9.8% of samples, and diverse vulnerability types covering multiple Common Weakness Enumeration (CWE) categories. To ensure fair comparison with baseline models, we used the identical data splits as reported in the original papers for each dataset.
For fine-tuning, we employed a dataset-specific approach to address the substantial class imbalance present in vulnerability detection. Class weighting was implemented to ensure balanced representation of vulnerable and safe code examples during training. Dataset-specific hyperparameter tuning included learning rates between 1e-5 and 5e-6, with smaller batch sizes (4) for fine-tuning to prevent overfitting.
Given the class imbalance in vulnerability detection, we prioritized metrics that provide a balanced assessment of model performance. F1 Score served as our primary metric, balancing precision and recall. We also evaluated using the Vulnerability Detection Score (VD-S) introduced by (Ding et al., 2024), which measures False Negative Rate. Accuracy, precision, and recall are reported for completeness and comparison with baseline models.
5 Experimental Results: Small Model, Big Impact
To evaluate White-Basilisk’s effectiveness, we conducted extensive experiments across five widely-used vulnerability detection datasets. Tables 1 and 2 present our results compared to state-of-the-art models. White-Basilisk consistently outperforms larger, more resource-intensive models across all benchmarks.
On the challenging PRIMEVUL dataset, White-Basilisk achieved an F1 Score of 0.2907 and VD-S of 0.7239, significantly outperforming larger models like StarCoder2 (7B parameters). Most impressively, on BigVul, our model achieved an F1 Score of 0.9490, accuracy of 0.9942, and VD-S of 0.0398, dramatically surpassing all competitors across all metrics. On the Draper dataset, White-Basilisk set a new standard with an F1 Score of 0.6069, while on REVEAL, it scored an F1 of 0.4934 and accuracy of 0.8988, outperforming previous best results. For VulDeepecker, our model achieved an F1 Score of 0.9388 with the highest precision at 0.9720.
Table 1: BigVul and PRIMEVUL Evaluation Results
Model PRIMEVUL BigVul
Acc F1 VD-S $\downarrow$ Acc F1 VD-S $\downarrow$
White-Basilisk 0.9630 0.2907 0.7239 0.9942 0.9490 0.0398
CodeT5 0.9667 0.1970 0.8993 0.9567 0.6493 0.7730
CodeBERT 0.9687 0.2086 0.8878 0.9557 0.6288 0.8177
UnixCoder 0.9686 0.2143 0.8921 0.9646 0.6546 0.6230
UnixCoder w/ balancing 0.9599 0.2628 0.8849 - - -
StarCoder2 0.9702 0.1805 0.8964 0.9620 0.6826 0.6914
CodeGen2.5 0.9665 0.1961 0.9151 0.9657 0.6730 0.6173
LineVul - - - - 0.9100 0.1400
Table 2: Draper, REVEAL, and VulDeepecker Evaluation Results
Model Draper REVEAL VulDeepecker
F1 Acc F1 F1 Prec
White-Basilisk 0.6069 0.8988 0.4934 0.9388 0.9720
Russell et al. (2018) 0.5660 - - - -
VulBERTa-MLP 0.4334 0.8448 0.4527 0.9303 0.9576
VulBERTa-CNN 0.5792 0.7973 0.4259 0.9086 0.9526
Baseline-BiLSTM 0.4684 0.7713 0.3911 0.6697 0.5258
Baseline-TextCNN 0.4940 0.7322 0.3741 0.7580 0.6348
REVEAL - 0.8437 0.4125 - -
VulDeepecker - - - 0.9290 0.9190
5.1 Ablation Studies: Validating Key Claims
To further validate our architectural choices and claims, we conducted three targeted ablation studies. Detailed results and analysis for each study are available in the Appendixes H, I, J.
5.1.1 Unified Model Analysis: Can One Model Rule Them All?
We trained a unified White-Basilisk model on the combined training data from all five datasets to assess cross-dataset generalization. This model maintained robust performance across diverse vulnerability types, with F1 scores of 0.938 on BigVul, 0.925 on VulDeepecker, 0.549 on Draper, 0.470 on REVEAL, and 0.233 on PRIMEVUL. Notably, the model maintained near-optimal performance on larger datasets while showing reasonable transfer to more challenging ones, demonstrating White-Basilisk’s ability to learn generalizable vulnerability patterns despite its compact size. The model achieved perfect detection (F1=1.000) for 22 different CWE categories in BigVul, including critical issues like out-of-bounds write (CWE-787) and cryptographic vulnerabilities (CWE-310).
Table 3: Performance on Key CWEs Across Datasets
CWE Description Datasets Samples F1 Range
CWE-119 Buffer Overflow BigVul, VulDeep, Draper 15,584 0.748-0.969
CWE-399 Resource Management BigVul, VulDeep 7,031 0.780-0.923
CWE-264 Access Control BigVul 1,240 0.925
CWE-476 NULL Pointer Deref. BigVul, Draper 1,709 0.686-0.971
CWE-416 Use After Free BigVul 963 0.958
5.1.2 Attention Efficiency Analysis: Linear vs. Quadratic Scaling
Does White-Basilisk’s linear-complexity attention truly enable efficient long sequence processing without performance degradation? We compared our approach against standard quadratic attention, measuring both memory usage and predictive performance across varying sequence lengths. As shown in Table 4, our approach achieves dramatic memory savings—using 24× less memory (1,338MB vs. 32,409MB) for sequences under 16K tokens while maintaining comparable or superior F1 scores (0.943 vs. 0.937). Crucially, while standard attention fails with out-of-memory errors beyond 16K tokens, White-Basilisk successfully processes sequences up to 131K tokens with only modest memory growth, demonstrating true linear scaling.
Most impressively, performance analysis revealed that White-Basilisk maintains strong detection capabilities even for extremely long sequences, with F1 scores of 0.943 for sequences under 16K tokens, 0.800 for 16K-32K tokens, and a perfect 1.000 for 32K-65K tokens. This confirms our architecture’s ability to capture long-range dependencies crucial for identifying vulnerabilities that span multiple functions or files. For detailed sequence-length analysis across all datasets, see Appendix I.
Table 4: Memory Usage and Performance by Sequence Length and Attention Type
Length Peak Memory (MB) BigVul F1 Score
Ours Standard Ours Standard
0-16K 1,338 32,409 0.943 0.937
16K-32K 1,654 OOM 0.800 -
32K-65K 2,213 OOM 1.000 -
65K-131K 3,322 OOM - -
OOM: Out of Memory on NVIDIA A100 40GB GPU
5.2 Efficiency and Environmental Impact
The remarkable performance of White-Basilisk becomes even more significant when considering its computational efficiency. At approximately 200M parameters, our model is orders of magnitude smaller than models like StarCoder2 (7B) while achieving superior results. This efficiency translates to substantial environmental benefits, as shown in Table 5.
Table 5: Comparison of CO2 Emissions
White-Basilisk CodeBERT StarCoder UnixCoder CodeT5 Gasoline Car (per Year) CO2 (kg) 85.5 2,240 23,000,000 2,048 1,136 4,600
We estimated the CO2 emissions during training using the Machine Learning Impact calculator (Lacoste et al., 2019). The environmental impact of White-Basilisk (85.5kg CO2) represents just 3.8% of CodeBERT’s emissions (2,240kg) and a mere 0.0004% of StarCoder’s emissions (23,000,000kg). This stark contrast underscores the environmental benefits of our approach, with CO2 savings equivalent to taking hundreds of cars off the road compared to larger model training.
The efficiency gains extend to inference as well. White-Basilisk can process sequences up to 128K tokens on a single NVIDIA A100 GPU, enabling whole-codebase analysis without the distributed computing infrastructure required by larger models. This practical efficiency makes advanced vulnerability detection accessible to a broader range of organizations and researchers.
Our results challenge the prevailing "bigger is better" paradigm in AI model development. White-Basilisk demonstrates that carefully designed architecture, rather than simply increasing parameter count, can yield superior performance while drastically reducing computational and environmental costs. This finding has significant implications for sustainable AI development, suggesting that the path forward may lie in architectural innovation rather than ever-larger models.
6 Limitations and Future Work
While White-Basilisk shows promising results in code vulnerability detection, it is important to acknowledge its current limitations and outline future directions. The main limitation of White-Basilisk is its focus on C and C++ codebases. The model’s ability to generalize across a broader range of programming languages, especially those with different syntaxes or paradigms, warrants further exploration. This constraint, combined with potential biases in our training and evaluation datasets, may limit the model’s generalizability to diverse real-world codebases. To address this, future work will involve expanding the model’s training to include a wider variety of programming languages and curating more representative datasets that reflect a broader spectrum of code samples and vulnerability types.
Also, despite strong performance metrics, White-Basilisk is not infallible. False positives and false negatives, particularly when detecting novel or zero-day vulnerabilities, remain an ongoing challenge. Additionally, while the model is capable of processing long sequences, its true understanding of complex, long-range dependencies in code still needs further investigation. Our future research will focus on reducing error rates, especially in high-stakes scenarios, and enhancing the model’s ability to analyze convoluted code structures spanning multiple functions or files.
Another area for improvement is the model’s explainability. Currently, White-Basilisk’s decision-making process is not sufficiently transparent. Improving the model’s ability to provide clear, actionable explanations for detected vulnerabilities is essential for building trust and delivering meaningful insights to developers. Future work will explore methods to offer context-aware understanding of the potential impact of vulnerabilities, as well as suggested fixes, ultimately aiming to evolve White-Basilisk into a comprehensive code analysis assistant rather than a mere detection tool.
White-Basilisk’s performance on a relatively small training dataset (2M samples) is impressive, but raises questions about potential limitations in its knowledge base compared to models trained on much larger datasets. To address this, we plan to scale White-Basilisk to approximately 1 billion parameters. While still modest in comparison to larger language models, this increase in parameter count aims to significantly boost performance while continuing to challenge the notion that only the largest models can deliver state-of-the-art results.
Beyond code vulnerability detection, there is potential for White-Basilisk’s architecture and training approach in broader AI applications. Future research will investigate its efficacy in various NLP tasks, exploring whether its computational efficiency and long-context capabilities can offer more resource-efficient alternatives to existing LLMs.
7 Conclusion
White-Basilisk represents a significant advancement in efficient AI model design, demonstrating that carefully architected compact models can outperform much larger alternatives in specialized domains. By achieving state-of-the-art results in code vulnerability detection with only 200M parameters, our work challenges the prevailing "bigger is better" paradigm in AI development. The model’s impressive performance—particularly its ability to process sequences up to 128,000 tokens—proves that context handling limitations can be addressed through architectural innovation rather than simply increasing model size.
The model’s extended context window addresses a fundamental limitation of many current LLMs, which often struggle with long-range dependencies and global code structure understanding. By enabling the analysis of entire codebases in a single pass, White-Basilisk can capture complex interdependencies and identify vulnerabilities that span multiple functions or files, a capability that has long eluded traditional approaches.
While the context-handling capabilities of White-Basilisk are its standout feature, it’s worth noting that these achievements have been realized with a relatively compact model of 200M parameters. This efficiency demonstrates that advances in AI are not solely dependent on increasing model size, but can also stem from innovative architecture design and training methodologies.
The implications of White-Basilisk’s approach extend beyond code vulnerability detection. The ability to handle extended contexts efficiently could prove valuable in numerous domains where long-range understanding is crucial, such as document analysis, complex system modeling, or long-form text generation. Moreover, the model’s efficiency opens up possibilities for deployment in resource-constrained environments, potentially bringing advanced AI capabilities to a broader range of applications and users.
In conclusion, White-Basilisk represents a significant step forward in addressing the context limitations of current LLMs, while also demonstrating that such advances need not come at the cost of excessive model size or computational requirements. As we continue to refine and expand upon this approach, we anticipate exciting developments in the field of AI, particularly in tasks that require deep understanding of extended contexts. The potential implications of this research are substantial, and we look forward to seeing how these ideas evolve and find application in diverse areas of AI and beyond.
Appendix A Appendix
Appendix B Evaluation Metrics
This appendix provides detailed descriptions of all metrics used to evaluate model performance in vulnerability detection tasks.
B.0.1 Accuracy
Accuracy measures the proportion of correct predictions (both true positives and true negatives) among all predictions:
$\text{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$ (7)
where TP = True Positives, TN = True Negatives, FP = False Positives, and FN = False Negatives.
B.0.2 Precision
Precision measures the proportion of correct positive predictions among all positive predictions:
$\text{Precision}=\frac{TP}{TP+FP}$ (8)
B.0.3 Recall
Recall (also known as sensitivity) measures the proportion of actual positives correctly identified:
$\text{Recall}=\frac{TP}{TP+FN}$ (9)
B.0.4 F1 Score
F1 Score is the harmonic mean of precision and recall, providing a balanced measure of model performance:
[…]
B.0.5 Vulnerability Detection Score (VD-S)
VD-S evaluates the False Negative Rate of a detector at a specific False Positive Rate (FPR) threshold:
[…]
where a lower score indicates better performance. This metric is particularly important for security applications as it measures the model’s ability to minimize missed vulnerabilities while maintaining a low false positive rate.
Each metric serves a specific purpose in evaluating different aspects of model performance, from general classification accuracy to specialized vulnerability detection capabilities. The combination of these metrics provides a comprehensive assessment of a model’s effectiveness in real-world security applications.
Appendix C Dataset Statistics Analysis
This section provides a comprehensive analysis of five vulnerability detection datasets (PRIMEVUL, BigVul, REVEAL, Draper, and VulDeepecker), examining their size distributions, class imbalance characteristics, and data quality metrics.
Table 6: Dataset Distribution and Vulnerability Statistics
Dataset Sample Count Vulnerable (%) Duplicates (%)
Train Val Test Train Val Test Train Val Test
Draper 1,019,471 127,476 127,419 6.46 6.47 6.48 0.00 0.00 0.00
PRIMEVUL 184,427 25,430 25,911 3.02 2.75 2.68 0.00 0.00 0.00
BigVul 150,908 18,864 18,864 5.79 5.88 5.59 0.005 0.00 0.00
VulDeepecker 128,118 16,015 16,015 6.08 6.08 6.08 37.72 20.27 20.33
REVEAL 18,187 2,273 2,274 9.90 9.24 10.11 1.17 0.18 0.09
Table 7: Sequence Length Statistics (Training Split)
Dataset Min Max Mean Median 95th %
PRIMEVUL 3 296,924 502.44 193.0 1,729.0
VulDeepecker 8 312,940 284.03 142.0 893.0
REVEAL 10 120,684 569.05 226.0 1,929.7
BigVul 6 70,440 343.90 147.0 1,193.0
Draper 10 42,492 320.85 236.0 858.0
C.1 Implications for Model Design
These statistics significantly influenced our model design decisions:
  1. 1.
    The substantial class imbalance across all datasets (ranging from 3.02% to 10.11% vulnerable samples) motivated our implementation of specialized class weighting and sampling strategies.
  2. 2.
    The extreme range in sequence lengths (from 3 to 312,940 tokens) justified our focus on developing an architecture capable of handling very long sequences efficiently.
  3. 3.
    The varying levels of data duplication (0% to 37.72%) highlighted the importance of robust evaluation metrics and careful interpretation of results, particularly for VulDeepecker.
  4. 4.
    The consistency of class distributions across splits suggests that our evaluation metrics should be reliable indicators of real-world performance.
  5. 5.
    REVEAL’s higher proportion of vulnerable samples (10%) compared to other datasets (3-6%) provides an important test case for our model’s ability to handle different class balance scenarios.
Appendix D Baseline Models
This section details the baseline models examined in our study. It is important to note that we did not train, finetune, or run any of these models ourselves. Instead, we collected and analyzed their reported metrics and configurations from their respective papers. For CodeT5 (CT5), CodeBERT (CB), UnixCoder (UC), StarCoder2 (SC2), and CodeGen2.5 (CG2.5), the information was sourced from the PrimeVul paper Ding et al. [2024]. For VulBERTa variants (VulBERTa-MLP and VulBERTa-CNN) and the baseline models (Baseline-BiLSTM and Baseline-TextCNN), the information was obtained from the original VulBERTa paper Hanif and Maffeis [2022]. For LineVul, the information was obtained from the original LineVul paper Fu and Tantithamthavorn [2022].
Table 8: Overview of baseline models examined in our study
Model Architecture Pre-training Params
CT5 Wang et al. [2021] Enc-Dec Multi-lingual code 220M
CB Feng et al. [2020] Encoder Bimodal (code + text) 125M
UC Guo et al. [2022] Encoder Cross-modal 125M
SC2 Li et al. [2023] Decoder The Stack v2 7B
CG2.5 Nijkamp et al. [2023] Decoder Code + natural lang. 7B
VulBERTa-MLP Hanif and Maffeis [2022] Encoder C/C++ code 125M
VulBERTa-CNN Encoder-CNN C/C++ code 2M
Baseline-BiLSTM BiLSTM None 1M
Baseline-TextCNN TextCNN None 1M
LineVul Fu and Tantithamthavorn [2022] BERT CodeBERT 125M
CT5: CodeT5, CB: CodeBERT, UC: UnixCoder,
SC2: StarCoder2, CG2.5: CodeGen2.5
Table 9: Training configurations as reported in respective papers
Configuration Value
Small Model Epochs ($<$7B) 10
Large Model Epochs (7B) 4
VulBERTa Pre-training Steps 500,000
VulBERTa Fine-tuning Epochs 10
BiLSTM/TextCNN Epochs 10
LineVul Fine-tuning Epochs 10
Appendix E Classification Head
The classification head of White-Basilisk is designed to efficiently transform the high-dimensional representations learned by the main model into classification outputs for vulnerability detection. Its architecture is as follows:
  1. 1.
    Dense Layer 1: A fully connected layer that projects the hidden state (dimension 512) to the same dimension. This layer uses a GELU activation function and is followed by dropout for regularization.
  2. 2.
    Dense Layer 2: Another fully connected layer that reduces the dimension from 512 to 256, again followed by GELU activation and dropout.
  3. 3.
    Layer Normalization: Applied to the output of Dense Layer 2 for improved stability and faster convergence.
  4. 4.
    Output Layer: A final linear layer that projects from 256 dimensions to the number of classes (typically 2 for binary classification of vulnerable vs. non-vulnerable code).
This classification head structure was chosen to gradually reduce the dimensionality of the representations while maintaining the model’s ability to capture complex patterns relevant to vulnerability detection. The use of GELU activations and layer normalization aligns with modern best practices in deep learning architecture design. The classification head is mathematically described as follows:
$x_{1}$ $=\text{Dropout}(\text{GELU}(W_{1}h+b_{1}))$ (12)
$x_{2}$ $=\text{Dropout}(\text{GELU}(W_{2}x_{1}+b_{2}))$
$x_{3}$ $=\text{LayerNorm}(x_{2})$
$y$ $=W_{3}x_{3}+b_{3}$
where $h\in\mathbb{R}^{512}$ is the input from the main model, $W_{1}\in\mathbb{R}^{512\times 512}$, $W_{2}\in\mathbb{R}^{256\times 512}$, and $W_{3}\in\mathbb{R}^{2\times 256}$ are learnable weights, and $b_{1},b_{2},b_{3}$ are biases.
Appendix F Hyperparameter Details
This section provides a detailed overview of the hyperparameters used in training White-Basilisk, including both the pretraining and fine-tuning phases. We also discuss the rationale behind key hyperparameter choices and their impact on model performance.
F.1 Pretraining Hyperparameters
Learning Rate: We chose a relatively small learning rate of 1.41e-5 to ensure stable training given the complexity of the task and the hybrid nature of our model architecture. This value was determined through careful tuning to balance training speed and convergence stability.
Batch Size: A batch size of 16 was selected as a compromise between training efficiency and memory constraints of our hardware (single NVIDIA A100 40GB GPU). Larger batch sizes could potentially improve training stability but would require more memory or gradient accumulation steps.
Number of Epochs and Warmup Ratio: We trained for 10 epochs with a warmup ratio of 0.15. This combination allowed the model to reach good performance while preventing overfitting. The warmup period helps stabilize training in the early stages.
Optimizer Settings: We used the AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1e-8$. These are standard settings that work well across a wide range of tasks. The weight decay of 0.01 was applied to all parameters except for bias and LayerNorm weights to prevent overfitting.
FIM and FIM-SPM Rates: Both the Fill-in-the-Middle (FIM) rate and the FIM Sentence Permutation Mode (SPM) rate were set to 0.5. This means that 50% of the samples undergo FIM transformation, and among those, 50% use the SPM variant. These rates provide a good balance between standard causal language modeling and the FIM objective, enhancing the model’s bidirectional understanding capabilities.
F.2 Fine-tuning Hyperparameters
Learning Rate and Batch Size: We used a smaller learning rate (5e-6) and batch size (4) for fine-tuning to prevent catastrophic forgetting and to allow the model to adapt to the specific characteristics of each dataset without overfitting.
SIFT Parameters: For Scale-Invariant Fine-Tuning, we used a learning rate of 1e-4 for the perturbation layer and an initial perturbation magnitude of 1e-2. These values were chosen to provide meaningful adversarial examples without overly distorting the input embeddings.
Appendix G Handling Class Imbalance
A significant challenge in the development of AI classification models is the management of highly imbalanced datasets. In such scenarios, it is important to train the model to maintain its ability to detect minority classes effectively. In the context of source code vulnerability detection, we also encounter highly imbalanced classification data. To address this issue, we employ a dual approach affecting both data sampling and loss computation. This is implemented via a weighted function $w_{c}$, where $w_{c}$ represents the weight for class $c$, and $N_{c}$ denotes the number of samples in class $c$. The function is defined as follows:
$w_{c}=\frac{N}{2N_{c}}$ (13)
Based on the provided function, we implement two key components in our methodology. Firstly, we employ a Weighted Random Batch Sampler, a sampling mechanism that ensures each mini-batch contains a balanced representation of classes, thus mitigating the effects of dataset imbalance during training. Secondly, we implement a weighted loss function, modifying the standard cross-entropy loss by incorporating the class-specific weights $w_{c}$.
G.0.1 SIFT (Scale-Invariant Fine-Tuning)
We implement automated adversarial training using SIFT to improve the model’s resilience against adversarial examples. SIFT operates by introducing small perturbations to the input during training, encouraging the model to learn more robust features. In our implementation, we added a PerturbationLayer into the model architecture, which applies learnable perturbations to the input embeddings. The training process was designed to minimize both the task loss and the adversarial loss, the latter being computed as the difference between predictions on clean and perturbed inputs.
This approach confers several advantages, including improved model generalization and enhanced robustness to minor variations in input. Such resilience is crucial in the domain of code vulnerability detection, where the model must maintain consistent performance across diverse code samples and potential adversarial inputs.
By adopting this technique, we have effectively imbued the model with a form of ’adversarial immunity’, rendering it more resilient against potential attacks or attempts to deceive its analysis. This enhanced robustness is particularly valuable in security-critical applications, where the reliability and consistency of the model’s performance are of paramount importance.
Appendix H Ablation Study: Combined Dataset Training
To further evaluate White-Basilisk’s performance and investigate the impact of training data composition, we conducted an ablation study using a combined training approach across all datasets. This experiment involved concatenating the training splits from all five datasets (REVEAL, Draper, VulDeepecker, BigVul, and PRIMEVUL) into a single unified training set. During each training epoch, the model was evaluated on the concatenated testing splits from all datasets to monitor for potential overfitting. Final performance metrics were obtained by evaluating the trained model separately on each dataset’s designated validation split, ensuring fair comparison with previous results.
H.1 Experimental Setup
The model was trained using the same hyperparameters as described in Section 6.2, but with the following data configuration:
  • Training Data: Combined training splits from all five datasets into a single training set
  • Testing: Concatenated testing splits from all datasets, used for monitoring training progress
  • Validation: Individual Validation splits for each dataset, evaluated separately to assess dataset-specific performance
This unified training approach resulted in a significantly larger and more diverse training set, allowing us to investigate how the model performs when exposed to a broader range of vulnerability patterns and coding styles simultaneously.
H.2 Results and Analysis
The results of this combined training approach are presented in Table 10.
Table 10: Combined Training Results Across All Datasets
Dataset Precision Recall F1 Accuracy
REVEAL 0.416 0.551 0.470 0.888
Draper 0.568 0.532 0.549 0.948
VulDeepecker 0.939 0.912 0.925 0.989
BigVul 0.936 0.940 0.938 0.993
PRIMEVUL 0.268 0.265 0.233 0.952
Combined Test 0.643 0.585 0.613 0.956
The model maintains strong performance across most datasets, with particularly robust results on VulDeepecker (F1: 0.925) and BigVul (F1: 0.938), suggesting effective transfer learning across different vulnerability detection tasks. Performance varies significantly across datasets, from an F1 score of 0.938 on BigVul to 0.233 on PRIMEVUL, indicating that some vulnerability patterns may be more challenging to learn in a combined setting. The model maintains high accuracy across all datasets (0.888-0.993), demonstrating robust overall classification performance even with the increased complexity of the combined training task. The model generally maintains a good balance between precision and recall, with some datasets showing nearly identical values (e.g., BigVul: 0.936/0.940), suggesting stable learning of vulnerability patterns.
H.3 Model Robustness Analysis
A particularly noteworthy aspect of these results is White-Basilisk’s ability to maintain stable performance across multiple diverse datasets without experiencing catastrophic forgetting or overfitting. This is especially significant given the model’s relatively compact size of 200M parameters. Several factors contribute to this robustness. The model shows signs of positive transfer learning, where knowledge gained from one dataset appears to benefit the detection of vulnerabilities in others. This is evidenced by the maintenance of high accuracy scores across all datasets despite their varying characteristics. Additionally, the model maintains performance across datasets of different sizes and complexity levels, from the smaller REVEAL dataset to the larger PRIMEVUL dataset. This stability suggests that the model’s learning capacity is well-matched to the task complexity.
H.4 Comparison with Individual Training
When compared to the individual training results presented in Section 5, the combined training approach shows some interesting trade-offs. For some datasets (VulDeepecker, BigVul), the performance remains close to individual training results, suggesting that the model can effectively learn and maintain dataset-specific patterns even in a combined setting. Performance on more challenging datasets like PRIMEVUL shows some degradation, indicating that the increased diversity of the training data may make it harder for the model to capture some of the more nuanced vulnerability patterns specific to certain datasets. The overall combined test metrics (F1: 0.613, Accuracy: 0.956) demonstrate that White-Basilisk can effectively learn from multiple datasets simultaneously while maintaining reasonable performance across all of them.
This ability to maintain stable performance across diverse datasets without overfitting or experiencing catastrophic forgetting is particularly notable for a model of this size. It suggests that White-Basilisk’s architecture strikes an effective balance between model capacity and efficiency, enabling robust multi-task learning without requiring the massive parameter counts typically associated with such capabilities. This finding has important implications for the development of efficient, multi-purpose vulnerability detection systems that can be deployed in resource-constrained environments while maintaining high performance across a range of vulnerability types.
Appendix I Ablation Study: Attention Mechanisms and Long-Range Vulnerability Detection
To thoroughly evaluate White-Basilisk’s performance and validate our architectural choices, we conducted a comprehensive ablation study focusing on two key aspects: (1) the effectiveness of our linear-complexity Infini-attention mechanism compared to standard self-attention, and (2) the model’s performance across varying sequence lengths. This analysis is particularly important given the prevalence of vulnerabilities that span multiple functions or files, requiring models to maintain effectiveness over long code sequences.
I.1 Experimental Setup
Model: We used the White-Basilisk checkpoint that was trained on the combined datasets from H.
We categorized sequences into four length bins for analysis:
  • Bin 1: 0-16,384 tokens (standard context length)
  • Bin 2: 16,384-32,768 tokens (extended context)
  • Bin 3: 32,768-65,536 tokens (long context)
  • Bin 4: 65,536-131,072 tokens (very long context)
For each bin, we monitored both memory consumption and model performance across all datasets, comparing our Infini-attention implementation against standard self-attention (eager implementation).
I.2 Memory Efficiency Analysis
Table 11 presents the peak memory consumption across different sequence lengths and attention implementations.
Table 11: Peak Memory Consumption (MB) by Sequence Length and Attention Type
Length Bin Infini-attention Standard Attention
Peak Reserved Peak Reserved
Bin 1 (0-16K) 1,338 1,442 32,409 39,696
Bin 2 (16K-32K) 1,654 1,956 OOM OOM
Bin 3 (32K-65K) 2,213 2,638 OOM OOM
Bin 4 (65K-131K) 3,322 3,848 OOM OOM
OOM: Out of Memory on NVIDIA A100 40GB GPU
The results demonstrate the significant memory advantages of our Infini-attention approach. Infini-attention shows near-linear memory scaling, increasing from 1.3GB to 3.3GB across bins. Standard attention requires 24x more memory for Bin 1 and fails entirely on longer sequences. While standard attention becomes infeasible beyond 16K tokens, Infini-attention successfully processes sequences up to 131K tokens with modest memory requirements.
I.3 Performance Analysis Across Sequence Lengths
Table 12 presents a comprehensive analysis of performance across different sequence lengths and datasets.
Table 12: Performance and Distribution Analysis Across Sequence Lengths
Sample Distribution Infini-attention Eager-attention
Dataset Bin Total Non-Vuln Vuln F1 Precision Accuracy F1 Precision Accuracy
BigVul Bin 1 18,853 17,747 1,106 0.943 0.946 0.993 0.937 0.967 0.993
Bin 2 7 5 2 0.800 0.667 0.857 - - -
Bin 3 4 3 1 1.000 1.000 1.000 - - -
VulDeePecker Bin 1 12,764 11,814 950 0.925 0.939 0.989 0.932 0.968 0.990
Bin 2 2 2 0 - - 1.000 - - -
Bin 3 2 2 0 - - 1.000 - - -
Bin 4 1 1 0 - - 1.000 - - -
PRIMEVUL Bin 1 25,411 24,715 696 0.233 0.208 0.952 0.240 0.237 0.958
Bin 2 18 15 3 0.250 0.200 0.667 - - -
Bin 3 1 1 0 - - 1.000 - - -
REVEAL Bin 1 2,269 2,062 207 0.474 0.416 0.888 0.468 0.445 0.898
Draper Bin 1 12,769 11,819 950 0.568 0.609 0.948 0.511 0.646 0.948
Several remarkable findings emerge from this analysis. White-Basilisk demonstrates remarkable effectiveness in detecting vulnerabilities in longer sequences (over 16K tokens), despite their relative rarity. In BigVul, the model achieves perfect detection (F1=1.000) for sequences in Bin 3. For Bin 2 sequences, it maintains strong performance (F1=0.800) despite the increased complexity. This effectiveness on longer sequences is particularly noteworthy given the increased difficulty of maintaining coherent attention over such distances. The infini-attention variant achieves comparable performance to full self-attention, with nearly identical metrics across major datasets (e.g., BigVul: 0.937 vs 0.943 F1). It maintains high precision while reducing computational complexity and demonstrates that linear attention is a viable alternative for vulnerability detection. In Bin 1, where most vulnerabilities occur, the model shows exceptional performance, with BigVul reaching F1=0.943, Accuracy=0.993 and VulDeePecker achieving F1=0.925, Accuracy=0.989. The model maintains effectiveness despite significant class imbalance, successfully detecting vulnerabilities even when they comprise only 2.74% of samples (PRIMEVUL) and maintaining balanced precision-recall trade-offs across length bins. Performance variations across datasets reveal interesting patterns. Stronger performance on BigVul and VulDeePecker suggests better handling of certain vulnerability types, while lower scores on PRIMEVUL indicate the challenge of detecting more subtle or complex vulnerabilities.
I.4 Comparative Analysis with Standard Attention
When comparing Infini-attention with standard attention (where possible), we observe. Infini-attention achieves comparable or better performance while using significantly less memory. Unlike standard attention, Infini-attention can process the full range of sequence lengths present in real codebases. The ability to handle longer sequences without performance degradation makes White-Basilisk suitable for analyzing entire codebases in a single pass.
I.5 Implications for Vulnerability Detection
This ablation study yields several important insights. The success of Infini-attention in maintaining high performance across sequence lengths validates our architectural choices. The model’s ability to handle sequences up to 131K tokens while maintaining accuracy demonstrates its practical utility for real-world applications. Strong performance on longer sequences suggests effective capture of long-range dependencies, crucial for detecting vulnerabilities that span multiple functions or files. The memory efficiency of our approach makes it feasible to deploy White-Basilisk on standard hardware, even for processing very long sequences.
These findings confirm that White-Basilisk successfully addresses the key challenges in vulnerability detection: maintaining high accuracy across varying sequence lengths while remaining computationally efficient. The model’s particular strength in handling long sequences, combined with its consistent performance on more common shorter sequences, makes it a practical and effective tool for real-world code security applications.
Appendix J Ablation Study: CWE-Specific Performance Analysis
To provide deeper insights into White-Basilisk’s vulnerability detection capabilities, we conducted a comprehensive analysis of its performance across different Common Weakness Enumeration (CWE) categories. This analysis focuses on the BigVul, Vuldeepecker and Draper dataset, which provide detailed CWE-level metrics, allowing us to understand the model’s strengths and limitations across various vulnerability types.
J.1 Experimental Setup
  • Model: We used the White-Basilisk checkpoint that was trained on the combined datasets from H.
  • Evaluation Splits: Validation split of each dataset
J.1.1 Dataset-Specific Performance Patterns
Draper Dataset Performance
In the Draper dataset (Table 14), we observe a consistent pattern of high precision (1.000) across all CWE categories, but with varying recall rates. CWE-119 (Buffer Overflow) achieves the highest recall (0.597) and F1 score (0.748). CWE-120 (Buffer Copy without Checking Size) shows similar performance (recall=0.592, F1=0.744). CWE-469 and CWE-476 demonstrate progressively lower recall (0.563 and 0.522 respectively). This pattern suggests that while the model is highly precise in its predictions, it exhibits some conservatism in vulnerability detection, particularly for less frequent vulnerability types.
VulDeePecker Dataset Analysis
The VulDeePecker results (Table 15) show more balanced precision-recall characteristics. CWE-119 demonstrates near-perfect balance (precision=0.939, recall=0.940). CWE-399 (Resource Management Errors) shows lower but consistent performance (precision=0.776, recall=0.785). The balanced metrics suggest more robust learning of these vulnerability patterns, possibly due to better representation in the training data.
BigVul Dataset Insights
The BigVul dataset (Table 16) provides the most comprehensive view of White-Basilisk’s capabilities across 50+ CWE categories. Several significant patterns emerge.
Regarding perfect detection cases, 22 CWE categories achieve perfect scores (F1=1.000). These include critical vulnerabilities such as CWE-787 (Out-of-bounds Write) and CWE-310 (Cryptographic Issues); access control issues including CWE-732 (Permission Assignment) and CWE-284 (Access Control); and various severity levels ranging from CWE-59 (Link Following) to CWE-617 (Reachable Assertion). It is notable that perfect detection spans both frequent (over 200 samples) and rare (under 50 samples) categories.
For high-volume vulnerability performance, CWE-119 (2,746 samples) shows excellent performance (F1=0.969), CWE-264 (1,240 samples) demonstrates strong results (F1=0.925), and CWE-20 (1,977 samples) exhibits robust detection (F1=0.933).
Performance degradation patterns are observed in certain vulnerability types. Resource-related vulnerabilities show more variable performance, with CWE-404 (Resource Shutdown) having the lowest F1 score (0.571) and CWE-772 (Missing Release) demonstrating lower precision (0.750). Format-string vulnerabilities (CWE-134) show precision-recall imbalance (0.500/1.000).
The impact of sample size reveals that large sample categories (over 1000 samples) show consistently strong but not perfect performance. Medium-sized categories (100-1000 samples) demonstrate more variable results. Small categories (under 100 samples) often show perfect or near-perfect scores, suggesting potential overfitting.
J.1.2 Cross-Dataset Performance Analysis
The model’s behavior across datasets reveals important patterns. Regarding CWE-119 consistency, as the only vulnerability type present across all three datasets, it shows interesting variation: BigVul with F1=0.969 (balanced precision-recall), VulDeePecker with F1=0.940 (balanced precision-recall), and Draper with F1=0.748 (high precision, lower recall). This variation suggests dataset-specific characteristics affect detection performance. For scale effects, larger datasets (BigVul) generally show more balanced precision-recall trade-offs compared to smaller datasets.
J.1.3 Implications and Insights
These results yield several important insights for vulnerability detection. Regarding architecture effectiveness, White-Basilisk’s strong performance across numerous CWE categories validates its hybrid architecture design for vulnerability detection.
Detection patterns indicate that memory-related vulnerabilities consistently show strong detection rates, resource management vulnerabilities present more challenges, and access control vulnerabilities demonstrate surprisingly robust detection.
Practical implications of these findings include that high precision across most categories suggests low false positive rates, variable recall in some categories indicates potential for missed vulnerabilities, and perfect detection in rare categories warrants further investigation for potential overfitting.
These findings demonstrate White-Basilisk’s strong general capability while highlighting specific areas for potential improvement. The comprehensive nature of these results, particularly in the BigVul dataset, provides strong evidence for the model’s practical utility in real-world vulnerability detection scenarios.
CWE Description BigVul Draper VulDeePecker
Samples F1 Samples F1 Samples F1
CWE-119 Buffer Overflow: Classic buffer overflow vulnerabilities 2,746 0.969 2,419 0.748 10,419 0.940
CWE-399 Resource Management Errors: Failures in managing system resources 1,435 0.923 - - 5,596 0.780
CWE-20 Input Validation: Improper input validation 1,977 0.933 - - - -
CWE-264 Access Control: Permissions, privileges, and access controls 1,240 0.925 - - - -
CWE-120 Buffer Copy: Buffer copy without checking size of input - - 4,750 0.744 - -
CWE-476 NULL Pointer Dereference 501 0.971 1,208 0.686 - -
CWE-416 Use After Free: Using memory after it has been freed 963 0.958 - - - -
CWE-200 Information Exposure: Exposure of sensitive information 883 0.944 - - - -
Table 13: Comparison of Most Frequent CWEs Across Datasets
CWE Precision Recall F1 Accuracy Total Pos. Ratio Neg. Ratio
CWE-119 1.000 0.597 0.748 0.597 2,419 1.000 0.000
CWE-120 1.000 0.592 0.744 0.592 4,750 1.000 0.000
CWE-469 1.000 0.563 0.721 0.563 252 1.000 0.000
CWE-476 1.000 0.522 0.686 0.522 1,208 1.000 0.000
CWE-other 1.000 0.472 0.642 0.472 3,579 1.000 0.000
Overall 0.609 0.532 0.568 0.948 127,476 0.065 0.935
Table 14: Draper Dataset Metrics for All CWEs
CWE Precision Recall F1 Accuracy Total Pos. Ratio Neg. Ratio
CWE-119 0.939 0.940 0.940 0.991 10,419 0.077 0.923
CWE-399 0.776 0.785 0.780 0.986 5,596 0.031 0.969
Overall 0.910 0.913 0.911 0.989 16,015 0.061 0.939
Table 15: VulDeePecker Dataset Metrics for All CWEs
CWE Prec. Rec. F1 Acc. Total Pos. % Neg. %
CWE-119 0.978 0.960 0.969 0.995 2746 8.27 91.73
CWE-20 0.905 0.963 0.933 0.992 1977 5.51 94.49
CWE-399 0.957 0.892 0.923 0.992 1435 5.16 94.84
CWE-264 0.925 0.925 0.925 0.994 1240 4.27 95.73
CWE-416 0.944 0.971 0.958 0.997 963 3.63 96.37
CWE-200 0.913 0.977 0.944 0.994 883 4.87 95.13
CWE-125 0.984 0.984 0.984 0.997 794 7.68 92.32
CWE-189 0.921 0.946 0.933 0.993 695 5.32 94.68
CWE-362 0.939 0.861 0.899 0.988 592 6.08 93.92
CWE-476 0.944 1.000 0.971 0.998 501 3.39 96.61
Table 16: Top 10 CWEs for BigVul by Sample Size